# Kairos: Towards Adaptive and Generalizable Time Series Foundation Models

### v1

@@ Abstract
Time series foundation models (TSFMs) have emerged as a powerful paradigm for time series analysis, driven by large-scale pretraining on diverse data corpora. However, time series inherently exhibit heterogeneous information density over time, influenced by system states and signal complexity, presenting significant modeling challenges especially in a zero-shot scenario. Current TSFMs rely on non-adaptive processing pipelines that fail to capture this dynamic nature. For example, common tokenization strategies such as fixed-size patching enforce rigid observational granularity, limiting their ability to adapt to varying information densities. Similarly, conventional positional encodings impose a uniform temporal scale, making it difficult to model diverse periodicities and trends across series. To overcome these limitations, we propose Kairos, a flexible TSFM framework that integrates a dynamic patching tokenizer and an instance-adaptive positional embedding. Kairos adaptively selects tokenization granularity and tailors positional encodings to the unique characteristics of each time series instance. Trained on a large-scale Predictability-Stratified Time Series (PreSTS) corpus comprising over 300 billion time points and adopting a multi-patch prediction strategy in the inference stage, Kairos achieves superior performance with much fewer parameters on two common zero-shot benchmarks, GIFT-Eval and the Time-Series-Library benchmark, consistently outperforming established methods across diverse tasks. The project page is at https://foundation-model-research.github.io/Kairos.
@@ 1 Introduction
Time series forecasting is a core component of many real-world applications, including cloud services, traffic management, retail, finance, and energy (Box et al., 2015; Hyndman & Athanasopoulos, 2018). Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data. This challenge has long motivated the development of few-shot and zero-shot forecasting methods (Oreshkin et al., 2020). More recently, inspired by the success of large-scale foundation models (FMs) (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building general-purpose time series foundation models (TSFMs). These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across a wide range of downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series. As shown in Figure 1(b) and exemplified in Figure 1(c), the statistical information density measured via spectral entropy on sliding windows (see Appendix D) varies not only across datasets but also across samples within the same series. However, existing models typically impose a uniform and invariant scale across all series, thereby overlooking internal variability. This holds true whether they tokenize each time point individually (Ansari et al., 2024; Shi et al., 2024), tokenize using fixed-size patches (Das et al., 2024), or model temporal dependencies using standard positional encodings (Liu et al., 2025).
This non-adaptive paradigm gives rise to two major challenges. (i) Local information extraction: Current tokenization strategies enforce fixed granularity, requiring the model to observe whole series with the same temporal resolution. Such rigidity prevents adaptation to heterogeneous or fluctuating information density, both across different series and within a single sequence. For instance, models cannot zoom in on fine-grained details during sudden market shocks, nor can they abstract more efficiently when the data remain stable. (ii) Temporal relationship modeling: Conventional positional encodings impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality. As a result, TSFMs struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics. Figure 1(b) and 1(d) illustrate these challenges, highlighting both the variability of information density within time series and the rigidity of current methods.
Figure 1: (a) The trade-off between performance (normalized MASE) and the number of parameters on GIFT-Eval benchmark (Aksu et al., 2024) for existing TSFMs. Our Kairos achieves a superior performance at a comparable parameter scale. (b) (c) Significant variation exists in information density across and within different time series datasets. (d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos utilizes a Mixture of Patch Tokenization to address dynamic changes in information density.
To address these challenges, we propose Kairos, a foundation model designed to handle the complexities of heterogeneous time series data. The name Kairos is derived from its mythological significance as the Greek god of the “opportune moment”, reflecting its adaptability. For local information extraction, we introduce a Mixture-of-Size Dynamic Patching (MoS-DP) mechanism that adaptively models time series at multiple granularities by selecting patch sizes based on local information density. Inspired by the computationally free null experts in Mixture-of-Experts architectures (Zeng et al., 2024; Jin et al., 2024), we further incorporate null patch sizes, which allow the model to flexibly vary the number of active patches. This design moves beyond static patching strategies, enabling more expressive modeling of nuanced, time-varying dynamics. For temporal relationship modeling, we propose Instance-adaptive Rotary Position Embedding (IARoPE), an extension of RoPE (Su et al., 2024), tailored for diverse temporal dynamics. Unlike standard RoPE, which represents positional information with a fixed temporal scale, our IARoPE modulates these scales using spectral features extracted from each series, thereby generating positional encodings uniquely adapted to individual instances.
@@
In addition to our core architectural innovations, we further introduce two key optimizations in the inference and training phases, respectively. For inference, we propose a multi-patch prediction strategy that simultaneously forecasts multiple future patches. This approach enhances forecasting robustness by mitigating cumulative errors and offers greater flexibility for variable-length prediction horizons. To further enhance training, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy. By prioritizing more predictable sequences while maintaining broad coverage, PreSTS provides high-quality supervision that supports efficient model scaling. As demonstrated by the results on the GIFT-Eval benchmark (Aksu et al., 2024) in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters. This strong performance is consistently observed on Time-Series-Library (TSLib) (Wang et al., 2024b) benchmarks as well. In summary, our main contributions are as follows:
  • We propose Kairos, a novel foundation model specifically designed to handle dynamically changing time series data. Kairos demonstrates strong resilience to commonly observed data variability in real-world time series.
  • The two key architectural innovations introduced, i.e., MoS-DP and IARoPE, address the challenges of varying information density and instance-specific temporal dynamics, thereby enhancing the generalization capability of foundation models for time series.
  • Through extensive empirical evaluation, we demonstrate the superior performance of Kairos with significantly fewer parameters on zero-shot forecasting benchmarks, highlighting its flexibility and adaptability across a wide range of diverse and non-stationary time series datasets.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). The scarcity of domain-specific time series data has long motivated research into models with strong few-shot or zero-shot learning capabilities (Oreshkin et al., 2020; 2021). Recently, inspired by the strong generalization of large language models, researchers are devoting more effort to TSFMs leveraging vast data corpora. Some works attempt to directly transfer the sequence modeling capabilities of large language models to the time series domain (Gruver et al., 2023). Others leverage the scalability of Transformer (Vaswani et al., 2017) and directly train unified foundation models on extensive time series corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025). A few studies have also explored lightweight architectures, including Multi-Layer Perceptrons (MLPs) such as TTM (Ekambaram et al., 2024). However, these models often apply fixed and unified tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization. Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive position embedding to enhance its generalizability for diverse time series data.
Time Series Tokenization. Leveraging Transformer model backbone requires an effective mechanism for local information extraction, i.e., to transform raw data into sequential elements (tokens), discretely or continuously (Sennrich et al., 2016; Agarwal et al., 2025). The intuitive approach for time series tokenization is to map each time point to a token. However, models employ this approach (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) suffer from sensitivity to noise and computational inefficiency. To address this, Nie et al. (2023) introduces a fixed-size patching method that became a widely adopted paradigm in many subsequent TSFMs (Woo et al., 2024; Liu et al., 2025). However, the fixed-size patches failed to flexibly adapt to time series with varying information densities. While models like Pathformer (Chen et al., 2024) and ElasTST (Zhang et al., 2024) attempt to address these limitations by improving feature fusion strategies, they still simply partition the entire sequence into a few predefined patch sizes, failing to dynamically adjust to the varying information densities in single time series data. Furthermore, as they are designed for dataset-specific training rather than as foundation models, they lack cross-domain generalization capabilities. In contrast, Kairos improves upon the existing methods by dynamically selecting appropriate patch sizes for each segment of a time series, rather than relying on one or more fixed patch sizes for the entire sequence.
@@
Position Embedding. Due to the position-unaware characteristic of self-attention mechanism, Transformers relies on position embeddings (PE) to model temporal information. Most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024). By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series. Existing adaptive PE methods in NLP (Zheng et al., 2024; Lin et al., 2024) are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series. In the time series domain, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data. However, its adaptation uses per-dataset training rather than dynamic adjustment. To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the effectiveness of time series foundation models.
Figure 2: The architecture of KAIROS, which including (1) Mixture-of-Size Dynamic Patching (MoS-DP): This module adaptively tokenizes the time series by fusing features from multiple granularities. As detailed in the expanded view, a Dynamic Patch Router first selects active experts (each corresponding to a patch size) for a coarsest patch and determines the finest patch size $p_{k}$ for tokenization. The final embedding for each resulting finest patch is then created via a hierarchical fusion process: it is the weighted sum of outputs from all activated experts that correspond to the granularity of the finest patch itself or its ancestors. For instance, the diagram illustrates how the embedding for a finest patch of size $p_{1}$ (marked ①) aggregates information from its ancestors of size $p_{2}$. This results in a sequence of fused embeddings rich with multi-scale information that is fed into the encoder. (2) Instance-Adaptive RoPE (IARoPE): This module (left) adjusts positional encodings for the Transformer by modulating them based on the unique frequency profile of each input series.
@@ 3.1 Overall Framework
We present the forecasting problem formulation: given the historical observations of a time series $\bm{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$, the objective is to forecast the future values $\bm{x}_{T+1:T+H}=(x_{T+1},\ldots,x_{T+H})\in\mathbb{R}^{H}$, where $T$ is the lookback window and $H$ denotes the forecast horizon.

### v2

@@ Abstract
Inherent temporal heterogeneity, such as varying sampling densities and periodic structures, has posed substantial challenges in zero-shot generalization for Time Series Foundation Models (TSFMs). Existing TSFMs predominantly rely on massive parameterization to absorb such heterogeneity, as their static tokenization and positional encoding schemes entangle diverse temporal patterns into a fixed representation space, encouraging memorization rather than adaptation. To address this limitation, we propose Kairos, a flexible and parameter-efficient TSFM that decouples temporal heterogeneity from model capacity through a novel tokenization perspective. Kairos introduces a dynamic patching tokenizer and a mixture-of-size encoding that adapt observational granularity to local information density, enabling fine-grained temporal abstraction without increasing model width or depth. In addition, we design a multi-granularity positional embedding based on dynamic rotary encodings, which conditions on instance-level spectral features and temporal structure induced by dynamic patching tokenization, allowing robust modeling of diverse temporal dependencies. Trained on a novel Predictability-Stratified Time-Series (PreSTS) corpus, Kairos achieves superior zero-shot performance with substantially fewer parameters on two mainstream benchmarks, GIFT-Eval and Time-Series-Library. The project page is at https://foundation-model-research.github.io/Kairos.
Figure 1: (a) Comparison on the GIFT-Eval benchmark (Aksu et al., 2024) demonstrates that our Kairos achieves superior zero-shot forecasting performance (lower normalized MASE) while requiring significantly fewer parameters than existing TSFMs. (b) (c) Significant variation exists in information density (i.e., signal complexity) across and within different time series datasets. (d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos uses Mixture-of-Size Tokenization to dynamically adapt to information density, enabling the learning of generalizable rules rather than memorization.
@@ 1 Introduction
Time series forecasting is a core component of many real-world applications (Box et al., 2015; Wang et al., 2024b). Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data. This challenge has long motivated the development of few-shot and zero-shot forecasting methods (Oreshkin et al., 2020). More recently, inspired by the success of large-scale foundation models (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building Time Series Foundation Models (TSFMs) for forecasting. These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across various downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series, severely compromising their parameter efficiency. As shown in Figure 1(b)(c), the statistical information density (details in Appendix G) varies not only across datasets but also across segments within the same series. By applying rigid architectures (exemplified in Figure 1(d)) to this inherently heterogeneous data, models are forced to use massive parameters to memorize specific patterns rather than acquiring adaptive, generalizable rules. Unfortunately, this strategy diminishes the overall utility of the model’s capacity, rendering it less effective for real-world deployment.
This deficiency manifests primarily in two fundamental aspects of model design: (i) Representation learning: TSFMs have yet to adopt dynamic tokenization strategies analogous to those in Large Language Models (LLMs). While LLMs utilize subword tokenization (Gage, 1994) to efficiently compress semantic units of varying lengths, TSFMs predominantly persist with static patching, whether through point-wise tokenization (Ansari et al., 2024; Shi et al., 2024) or fixed-size patches (Das et al., 2024). By coupling token generation to fixed time intervals rather than information content, these approaches forcibly compress diverse temporal variations into a rigid representation space. Consequently, the model wastes computational capacity on low-information segments while struggling to capture high-frequency details within the same fixed budget, leading to significant parameter inefficiency. (ii) Temporal dependency modeling: Conventional positional encodings in TSFMs impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality. As a result, existing TSFMs like (Liu et al., 2025) struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics, as we empirically validate in Section 4.3. Consequently, the model requires a deeper, more complex architecture to compensate for this misalignment, further exacerbating parameter inefficiency and hindering practical utility.
To address these limitations, we propose Kairos, a parameter-efficient model comprising a Mixture-of-Size Encoder, a Heterogeneity-Aware Transformer, and a Multi-Patch Decoder. Kairos addresses time series heterogeneity through two primary mechanisms: (i) For intra-patch representation learning, the Mixture-of-Size Encoder adaptively models time series at multiple granularities based on local characteristics. Inspired by null experts in Mixture-of-Experts (Zeng et al., 2024; Jin et al., 2024), we introduce null patch sizes to dynamically adjust the number of active granularities. This approach moves beyond static tokenization strategies, enabling more expressive modeling of nuanced, time-varying dynamics. (ii) For inter-patch dependency modeling, we enhance the Transformer with Dynamic Rotary Position Embedding (DRoPE). Unlike standard RoPE (Su et al., 2024), which represents positional information with a fixed temporal scale, DRoPE modulates temporal scales using spectral features and granularity to generate instance-specific positional encodings. Flexible intra-patch and inter-patch modeling enable Kairos to effectively characterize heterogeneous time series with a substantially reduced parameter count (see Figure 1(a)). Additionally, the Multi-Patch Decoder uses learnable tokens to predict future patches in parallel, which mitigates cumulative errors in autoregressive generation and offers greater flexibility for variable-length prediction horizons.
@@
To further enhance training, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy. By prioritizing more predictable sequences while maintaining broad coverage, PreSTS provides high-quality supervision that supports efficient model scaling. As demonstrated by the results on the GIFT-Eval benchmark (Aksu et al., 2024) in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters. This strong performance is consistently observed on Time-Series-Library (TSLib) (Wang et al., 2024b) benchmarks as well.
In summary, our main contributions are as follows: (i) We present Kairos, a parameter-efficient time series foundation model designed to handle the dynamic and heterogeneous nature of time series data. (ii) We introduce three key architectural innovations: the Mixture-of-Size Encoder for adaptive multi-granularity tokenization, DRoPE for instance-specific temporal calibration, and the Multi-Patch Decoder for flexible and efficient parallel forecasting. (iii) Extensive empirical evaluations on the GIFT-Eval and Time-Series-Library benchmarks confirm the efficacy of Kairos, highlighting its superior zero-shot forecasting performance.
@@ 2 Related Work
Time Series Foundation Models (TSFMs). Inspired by the strong generalization of Large Language Models (LLMs), researchers have devoted increasing attention to TSFMs. Existing approaches typically involve adapting LLMs to time series (Gruver et al., 2023), training scalable Transformers on large corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025), or exploring alternative architectures, including non-Transformer models (Auer et al., 2025; Graf et al., 2025) and lightweight Multi-Layer Perceptrons (Ekambaram et al., 2024). However, these models often apply fixed tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization. Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive, granularity-aware position embedding to enhance its generalizability for diverse time series data.
Time Series Tokenization. Transformer-based models require converting raw data into discrete or continuous tokens (Sennrich et al., 2016; Agarwal et al., 2025). Current strategies generally fall into two static paradigms: (i) Fixed-size: The dominant paradigm relies on a uniform patch size across the sequence. This ranges from fine-grained point-wise tokenization (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) to coarser uni-size patching (Nie et al., 2023), which has been adopted by recent foundation models (Woo et al., 2024; Liu et al., 2025). While the latter improves upon the computational inefficiency of point-wise methods, both remain rigid. (ii) Multi-size: To capture broader temporal dynamics, models like Pathformer (Chen et al., 2024) and TTM (Ekambaram et al., 2024) incorporate multiple predefined patch sizes. However, these scales are applied statically, rather than adapting to the time series instance. All the above approaches partition sequences based on fixed time intervals rather than local information. This rigidity forces the model to allocate equal parameters to segments of varying complexity. This imbalance inevitably leads to suboptimal parameter efficiency and limited generalization. In contrast, Kairos employs a mixture-of-size strategy. By allocating fewer tokens to simple patterns and more to complex regions, Kairos significantly enhances parameter efficiency while maintaining robust generalization.
@@
Position Embedding. Due to the position-unaware characteristic of self-attention mechanism, Transformers Vaswani et al. (2017) rely on position embeddings (PEs) to model temporal information. However, most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024). By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series. Existing adaptive PE methods in NLP (Zheng et al., 2024; Lin et al., 2024) are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series. Therefore, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data. However, its adaptation uses per-dataset training rather than dynamic adjustment. To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the parameter efficiency and overall performance of TSFMs.
Figure 2: The architecture of Kairos, which includes (i) The Mixture-of-Size Encoder (right) adaptively tokenizes the input series by routing segments to optimal granularity experts and fusing their embeddings into a unified sequence $\boldsymbol{H}$. (ii) The Heterogeneity-Aware Transformer (middle) processes these tokens using Dynamic Rotary Position Embedding (DRoPE) (left), which modulates frequencies based on instance-level spectral features and calibrates positions to account for varying patch sizes. (iii) The Multi-Patch Decoder (top-middle) employs learnable forecast tokens to predict multiple future patches in parallel via cross-attention.
@@ 3.1 Overall Framework
We formulate our task as learning a time series forecasting function $f_{\phi}:\mathbb{R}^{T}\to\mathbb{R}^{H}$ parameterized by $\phi$. Given historical time series $\mathbf{X}\in\mathbb{R}^{T}$, time series forecasting is to predict the future horizon $\mathbf{Y}\in\mathbb{R}^{H}$. The learning objective is to minimize the discrepancy between the ground truth $\mathbf{Y}$ and the prediction $\hat{\mathbf{Y}}=f_{\phi}(\mathbf{X})$. We detail it in Appendix D.1.2.